### v1

@@ Abstract.
Let $X$ be a metric space with bounded geometry. We show that if $X$ admits a coarse embedding into an $\ell^{p}$-space ($1\leq p<\infty$), then the canonical inclusion from any geometric ideal to the corresponding ghostly ideal induces an isomorphism in $K$-theory. Our approach relies on the construction of twisted Roe algebras, avoiding the use of groupoid techniques. As consequences, we deduce the relative (maximal) coarse Baum-Connes conjectures for such spaces, as well as the Operator Norm Localization property for finite rank projections ($ONL_{\mathcal{P}_{Fin}}$).
@@ 1. Introduction
Ghosts are among the most mysterious operators in the Roe algebra. Let $X$ be a metric space with bounded geometry and fixed base point $x_{0}$. We denote $C^{*}(X)$ the Roe algebra of $X$ (2.2). An operator $T\in C^{*}(X)$ can be written as a $X$-by-$X$ matrix. Such $T$ is called a ghost if its matrix entries vanish at infinity; that is, $T_{xy}\to 0$ as $x,y\to\infty$. However, despite this local vanishing behavior, the operator $T$ is not necessarily compact. In other words, while the entries fade locally, the operator norm $\|\chi_{K_{R}}T\chi_{K_{R}}\|$ (where $K_{R}=B(x_{0},R)$) may not tend to zero as $R\to\infty$. Consequently, a ghost is an object that is elusive locally at infinity but remains capturable globally. The collection of all such operators forms a closed ideal in the Roe algebra, known as the ghostly ideal. It is also widely known as the ghost ideal in the literature. For a detailed discussion on the construction of ghost projections and the structural analysis of ghostly ideals, we refer the reader to [CW2004, WYI2012, KLVZ2021, LSZ2023, WZ2023].
Ghostly ideals play a crucial role in the study of the coarse Baum-Connes conjecture by providing counterexamples. Let $X$ be a metric space with bounded geometry. There is a coarse assembly map given by
@@
The coarse Baum-Connes conjecture states that this map is an isomorphism. However, if $X$ is a sequence of expander graphs, there exists a ghost projection in the Roe algebra $C^{*}(X)$. This projection defines an element in $K_{0}(C^{*}(X))$ that lies outside the image of the assembly map $\mu$, as shown in [HLS2002, WYI2012].
On the other hand, the study of ideal structures in Roe algebras is also fundamental for $K$-theoretic computations. In [HRY1993], N. Higson, J. Roe, and G. Yu initially introduced geometric ideals associated with subspaces. They utilized this structure to establish the coarse Mayer-Vietoris principle, which allows for the reduction of $K$-theory calculations to subspaces. In [CW2004], X. Chen and Q. Wang extended the concept of geometric ideals from subspaces to invariant open sets in the coarse groupoid (see 2.5). They proved that every ideal in the uniform Roe algebra corresponds to an invariant open subset of the coarse groupoid. Furthermore, they proved that if the metric space $X$ has Yu’s Property A, then every ideal in the uniform Roe algebra is geometric. Later in [RW2014], J. Roe and R. Willett proved the converse, establishing that every ideal of the uniform Roe algebra is geometric if and only if the space has Property A.
@@
More recently, Q. Wang and J. Zhang [WZ2023] introduced a refined concept of ghostly ideals associated with specific invariant open sets (see 2.5). Roughly speaking, for an invariant open set $U$, the corresponding ghostly ideal consists of operators whose matrix entries $T_{xy}$ vanish as $(x,y)$ tends to infinity along the direction of the complement $U^{c}$, rather than requiring them to vanish in all directions at infinity. When the subspace is bounded, the corresponding geometric ideal coincides exactly with the algebra of compact operators, and the corresponding ghostly ideal coincides with the original ghostly ideal of all ghost operators.
Although the requirement for these two ideals to be isomorphic at the algebraic level is very stringent (specifically, Property A), the conditions for their isomorphism at the $K$-theory level are much less restrictive. In [FS2014], M. Finn-Sell proved that if a space $X$ admits a coarse embedding into a Hilbert space, the inclusion map from the compact operator algebra to the ghost ideal induces an isomorphism on $K$-theory. In [WZ2023, WFZ2025], this result was generalized to the embedding from general geometric ideals to ghostly ideals.
The coarse Baum-Connes conjecture is closely related to coarse embeddings into Hilbert spaces. G. Yu’s milestone result proved that if a metric space with bounded geometry admits a coarse embedding into Hilbert space, then the coarse Baum-Connes conjecture holds for it. This condition implies the a-T-menability of the coarse groupoid $G(X)$, see [STY2002]. Indeed, previous results regarding the $K$-theory isomorphism between geometric and ghostly ideals rely heavily on the a-T-menability of the coarse groupoid.
@@
Recently, the connection between the coarse Baum-Connes conjecture and the geometry of $\ell^{p}$-spaces ($1\leq p<\infty$) has become increasingly tight. In [WXYZ2024], it proves that if a space coarsely embeds into an $\ell^{p}$-space, then the coarse Baum-Connes conjecture holds for $X$. Furthermore, it is shown that the coarse Novikov conjecture holds if $X$ admits a fibred coarse embedding into an $\ell^{p}$-space in [GLWZ2024]. A natural question arises: does the $K$-theory isomorphism between geometric and ghostly ideals persist for spaces coarsely embeddable into $\ell^{p}$-spaces? In this paper, we prove the following result:
Theorem 1.1 (2.6 & 3.16).
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space (or $L^{p}$-space) with $p\in[1,\infty)$, then for any invariant open subset $U$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism on $K$-theory, i.e.,
@@
is an isomorphism.
We emphasize that a-T-menability is linked to Hilbert space geometry. For a space $X$ that coarsely embeds into an $\ell^{p}$-space, its coarse groupoid is not necessarily a-T-menable. Consequently, our approach differs essentially from the methods used in [WZ2023, WFZ2025]. Inspired by [WXYZ2024], we employ the Dirac-dual-Dirac construction to reduce the problem to the study of ideals in twisted algebras. At the level of twisted algebras, the structure readily implies the $K$-theory isomorphism between these two ideals.
As applications of our main result, we derive the following corollaries. The first corollary concerns the relative coarse Baum-Connes conjecture, introduced in [GWZ2025]. Its validity characterizes local obstructions to positive scalar curvature at infinity for manifolds.
@@
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$). For any subspace $Y$, the relative coarse Baum-Connes conjecture holds for $(X,Y)$.
The second concerns the operator norm localization for equi-approximable finite-rank projections, a concept introduced in [BFV2024] that is of interest in geometric group theory. In this paper, we establish a connection between this condition and the geometry of $\ell^{p}$-spaces via $K$-theory computations.
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into $\ell^{p}$, then $X$ has operator norm localization for equi-approximable finite-rank projections.
Finally, we extend our results from the reduced setting to the maximal setting. By combining our main theorem with the techniques in [HIT2020, WXYZ2024], we deduce the following result regarding the maximal coarse Baum-Connes conjecture:
@@
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ induces an isomorphism on $K$-theory, i.e.,
@@
    For any subspace $Y$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ induces an isomorphism on the level of $K$-theory, i.e.,
@@
    For any subspace $Y$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
Our results naturally motivate further inquiries regarding more general coefficients and embedding conditions. We propose the following open questions:
@@
In [DG2024], a twisted version of the coarse Baum-Connes conjecture is introduced. It is natural to ask whether the twisted coarse Baum-Connes conjecture holds for $X$ with any twisted coefficient algebra if $X$ admits a coarse embedding into $\ell^{p}$-spaces?
Furthermore, parallel with [CWY2013], does the maximal coarse Baum-Connes conjecture hold for $X$ if $X$ admits a fibred coarse embedding into $\ell^{p}$-space, with $p\in[1,\infty)$?
The paper is organized as follows. In Section 2, we recall some preliminaries on coarse geometry, and ideal structures of Roe algebras. Section 3 is devoted to the proof of our main theorem in four steps. Finally, in Section 4, we discuss applications to the relative and maximal coarse Baum-Connes conjectures, as well as the operator norm localization property.
@@ 2.1. Coarse groupoid
where $B(x,R)=\{x^{\prime}\in X\mid d(x,x^{\prime})\leq R\}$. Denote $\beta X$ the Stone-Čech compactification of $X$, which is the set of all ultrafilters on $X$ endowed with a topology making it a compact Hausdorff space such that $C(\beta X)\cong\ell^{\infty}(X)$. The reader is referred to [Goldbring2022] for details of ultrafilters and Stone-Čech compactification. For each $R>0$, we define the $R$-diagonal of $X\times X$ by
@@
When $R=0$, the set $\Delta_{0}$ is exactly the diagonal in $X\times X$. A subset of an $R$-diagonal is called an entourage (or a controlled set). The set of all entourages, denoted by $\mathcal{E}_{d}$, is called the coarse structure of $X$ associated with the metric. The coarse groupoid of $X$, denoted by $G(X)$, is defined to be
@@
The space $\beta X$ is the unit space of $G(X)$. The projection onto the first (second, resp.) coordinate $r:X\times X\to X$, ($s:X\times X\to X$, resp.) extends continuously to the range map $r:G(X)\to\beta X$ (the source map $s:G(X)\to\beta X$, resp.). The reader is referred to [STY2002] for details of the coarse groupoid. A subset $U\subset\beta X$ is called invariant if $r(\gamma)\in U$ is equivalent to $s(\gamma)\in U$ for any $\gamma\in G(X)$. For a subset $Y\subseteq X$, we define
@@
It is proved in [SpaWil2017, Appendix C] that for any $\gamma\in G(X)$ with $s(\gamma)=\alpha\in\beta X$, there exists a partial translation $t:D\to R$ in $X$ such that $\gamma=\lim_{x\to\alpha}(t(x),x)\in G(X)$.
For any map $f:X\to Y$, we shall still denote $f$ the extended continuous map $\beta X\to\beta Y$ with a little abuse of notation. For any subset $A\subseteq\beta X$, we denote $G(X)|^{A}_{A}=r^{-1}(A)\cap s^{-1}(A)$, which is subgroupoid. Let $Y\subseteq X$ be a subset. Then the canonical inclusion map $i:Y\to X$ extends to $i:\beta Y\to\beta X$. Thus, one can identify $\beta Y$ to the closure of $Y$ in $\beta X$. As a result, we have that $G(Y)=G(X)|^{\beta Y}_{\beta Y}$. For any invariant subset $U\subseteq\beta X$, we shall denote $U^{\beta Y}=\beta Y\cap U$ and it is direct to see that $U^{\beta Y}$ is $G(Y)$-invariant. A coarse ideal pair $(X,U)$ includes a metric space $X$ and an invariant subset $U$. Two pairs $(X,U)$ and $(Y,V)$ are coarsely equivalent if there exists a coarse equivalence $f:X\to Y$ such that $V$ is the smallest invariant open set containing $f(U)$. In this case, it is equivalent to say that $V$ is the minimal invariant open subset such that $U=f^{-1}(V)$.

### v2

@@ Abstract.
Ghostly ideals are among the most mysterious objects in coarse index theory. In this paper we show that if a metric space $X$ with bounded geometry admits a coarse embedding into an $\ell^{p}$-space ($1\leq p<\infty$), then the canonical inclusion from any geometric ideal to the corresponding ghostly ideal induces an isomorphism in $K$-theory. As consequences, we deduce that such spaces satisfy the relative coarse Baum-Connes conjectures introduced in Guo et al. [2025], as well as the operator norm localization property for finite rank projections ($ONL_{\mathcal{P}_{Fin}}$) as introduced in Braga et al. [2024].
Key words and phrases:
Coarse geometry, Roe algebra, Ghost ideal, $K$-theory, $\ell^{p}$-spaces
2020 Mathematics Subject Classification:
19K56, 47L20
Corresponding author: liangguo@simis.cn
@@ 1. Introduction
The notion of ghosts was introduced by Guoliang Yu in his study of the coarse Baum-Connes conjecture and it plays a very central role in both coarse geometry and operator algebra. Ghosts are among the most mysterious operators in Roe algebras: a ghost is an infinite matrix such that its matrix entries vanish at infinity, yet it may fail to be a compact operator. More precisely, let $C^{*}(X)$ denote the Roe algebra of a metric space $X$ with bounded geometry, i.e., there is a uniform bound on the cardinalities of all closed balls of a fixed radius (see 2.2). An operator $T\in C^{*}(X)$ can be written as a $X$-by-$X$ matrix such that each matrix entry is a compact operator, and $T$ is called a ghost if $\|T_{xy}\|\to 0$ as $x,y\to\infty$. However, despite this local vanishing behavior of its matrix entries, the operator $T$ itself is not necessarily compact. In other words, the operator norm $\|\chi_{K_{R}}T\chi_{K_{R}}\|$ may not tend to zero as $R\to\infty$, where $K_{R}=B(x_{0},R)$ for any fixed base point $x_{0}\in X$ and $R>0$. Consequently, a ghost is an object that is elusive locally at infinity but remains capturable globally.
As mentioned previously, ghosts play a crucial role in the study of the coarse Baum-Connes conjecture by providing counterexamples. For a metric space $X$ with bounded geometry, there is a so-called coarse assembly map
@@
The coarse Baum-Connes conjecture states that the map $\mu$ is an isomorphism of abelian groups. However, if $X$ is a sequence of expander graphs, there exists a non-compact ghost projection in the Roe algebra $C^{*}(X)$. This projection defines an element in $K_{0}(C^{*}(X))$, which violates the conjecture because it lies outside the image of the assembly map $\mu$, as shown in Higson et al. [2002]; Willett and Yu [2012]. For generalized results extending from expander graphs to (measured) asymptotic expanders, see Khukhro et al. [2021]; Li et al. [2023]. On the other hand, the most significant sufficient condition for the coarse Baum-Connes conjecture was established in the milestone work of G. Yu in 2000, who proved that the conjecture holds for any metric space of bounded geometry admitting a coarse embedding into a Hilbert space (see Yu [2000]). This result established a crucial link between coarse geometry of Hilbert spaces and the analytic index theory.
In analogy with the closed ideal $\mathcal{K}$ of all compact operators, the collection of all ghosts in a Roe algebra forms a closed ideal $\mathcal{G}$ in the Roe algebra, referred to as the ghostly ideal [^1] It is also frequently referred to in the literature as the ghost ideal.. It turns out that the ghostly ideal $\mathcal{G}$ contains $\mathcal{K}$ and is in fact directly linked to the coarse Baum-Connes conjecture. Indeed, M. Finn-Sell in [Finn-Sell, 2014, Proposition 35] proved that the inclusion map from the compact ideal $\mathcal{K}$ into the ghostly ideal $\mathcal{G}$ induces an isomorphism in $K$-theory if and only if the coarse assembly map $\mu$ is an isomorphism, provided the space $X$ admits a fibred coarse embedding into a Hilbert space. The notion of fibred coarse embedding into Hilbert space first introduced by Chen, Wang and Yu in Chen et al. [2013], and it greatly generalizes Gromov’s notion of coarse embedding into Hilbert space. As consequences of Yu [2000]; Finn-Sell [2014], if a space $X$ coarsely embeds into a Hilbert space, the inclusion map from the compact ideal $\mathcal{K}$ into the ghostly ideal $\mathcal{G}$ induces an isomorphism in $K$-theory (see [Finn-Sell, 2014, Corollary 36]).
@@
In recent years, there has been an uptick in interest in the study of general geometric ideal $\mathcal{I}(X,U)$ and ghostly ideal $\mathcal{G}(X,U)$ associated with a specific invariant open subset $U$ in $\beta X$ (see Wang and Zhang [2025]; Wang et al. [2025]). Roughly speaking, for an invariant open set $U$ in $\beta X$, the corresponding ghostly ideal consists of operators whose matrix entries $T_{xy}$ vanish as $(x,y)$ tends to infinity along the ($\U\beta X\backslash$ U)-direction, rather than requiring matrix entries to vanish in all directions at infinity. We refer the reader to 2.5. When $U=X$, the corresponding geometric ideal $\mathcal{I}(X,X)$ coincides exactly with the compact ideal $\mathcal{K}$, and the corresponding ghostly ideal $\mathcal{G}(X,X)$ coincides with the original ghostly ideal $\mathcal{G}$ of all ghost operators in the Roe algebra. On the other hand, we recover the Roe algebra when $U=\beta X$ in the sense that $\mathcal{I}(X,\beta X)=\mathcal{G}(X,\beta X)=C^{*}(X)$. Moreover, it is known that all ghostly ideals in the Roe algebra are geometric if and only if the underlying space $X$ has Property A (see e.g. [Wang et al., 2025, Proposition 5.1]), and the canonical inclusion from any geometric ideal $\mathcal{I}(X,U)$ to the corresponding ghostly ideal $\mathcal{G}(X,U)$ induces an isomorphism in $K$-theory if the underlying space $X$ coarsely embeds into a Hilbert space (see [Wang et al., 2025, Theorem 5.3]).
@@
Progress on the coarse Baum-Connes conjecture remained limited until recently, when the connection between the coarse Baum-Connes conjecture and the coarse geometry of $\ell^{p}$-spaces ($1\leq p<\infty$) became clearer. The principal achievements in this direction establish the injectivity of the coarse assembly map $\mu$ for spaces admitting a fibred coarse embedding into an $\ell^{p}$-space (see Guo et al. [2024]), and the bijectivity of $\mu$ for spaces admitting a coarse embedding into an $\ell^{p}$-space (see Wang et al. [2024]). A natural question arises in this perspective: does the $K$-theory isomorphism between the geometric ideal $\mathcal{I}(X,U)$ and the ghostly ideal $\mathcal{G}(X,U)$ persist for the space $X$ that coarsely embeds into an $\ell^{p}$-space? In this paper, we answer this question affirmatively (even for spaces coarsely embeddable into an $L^{p}$-space):
Theorem 1.1 (2.6 & 3.17).
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space (or $L^{p}$-space) with $p\in[1,\infty)$, then for any invariant open subset $U\subseteq\beta X$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism in $K$-theory, i.e.,
@@
is an isomorphism of abelian groups, where $\mathcal{I}(X,U)$ and $\mathcal{G}(X,U)$ are the geometric and ghostly ideal associated with $U$ in the Roe algebra $C^{*}(X)$, respectively.
It should be emphasized that the proofs of the $K$-theory isomorphism between the geometric and ghostly ideals in Wang and Zhang [2025]; Wang et al. [2025] depend crucially on the fact that a-T-menable groupoids are $K$-amenable (see Tu [1999]). Indeed, a metric space $X$ with bounded geometry is coarsely embeddable into a Hilbert space if and only if the associated coarse groupoid $G(X)$ is a-T-menable (see [Skandalis et al., 2002, Theorem 5.4]). For a space $X$ coarsely embeddable into an $\ell^{p}$-space, whether its coarse groupoid $G(X)$ is $K$-amenable remains an open question. It is precisely this gap that necessitates a fundamentally different approach in this paper, departing from earlier methods.
Our method to prove 1.1 is to generalize the $\ell^{p}$-Dirac-dual-Dirac construction in Wang et al. [2024] from the extreme case where $U=\beta X$ to a general invariant open subset $U\subseteq\beta X$. The argument proceeds in three main stages:
1. Reduction to sparse subspaces: Since the $\ell^{p}$-Dirac-dual-Dirac construction is primarily applicable to sparse spaces, we first reduce the proof of 1.1 to the sparse subspaces of $X$ (see 2.12).
2. Restriction of the construction to ideals: On a sparse subspace, the $\ell^{p}$-Dirac-dual-Dirac construction restricts to the ghostly and geometric ideals, respectively (see 3.13). This allows us to further reduce the proof to a comparison between the geometric and ghostly ideals within a twisted Roe algebra.
3. Comparison within the twisted framework: In the twisted Roe algebra, the distinction between elements in the geometric ideal and the corresponding ghostly ideal vanishes asymptotically (see 3.14), which completes the proof of our main theorem.
Several corollaries follow as direct applications of our main result. The first corollary concerns the so-called relative coarse Baum-Connes conjecture, introduced in Guo et al. [2025]. Its validity provides a local obstruction to the existence of Riemannian metrics with positive scalar curvature at infinity (see Guo et al. [2025] for details).
@@
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$). For any subspace $Y$ in $X$, the relative coarse Baum-Connes conjecture holds for $(X,Y)$.
The second involves $\mathrm{ONL}_{\mathcal{P}_{\mathrm{Fin}}}$—the operator norm localization property for equi-approximable finite-rank projections introduced in Braga et al. [2024]. This property, which plays a role in rigidity problems and the structure of Cartan masas in uniform Roe algebras, is a weaker version of the classical operator norm localization property Chen et al. [2008]. It is known that $\mathrm{ONL}_{\mathcal{P}_{\mathrm{Fin}}}$ is equivalent to the condition that all ghost projections on every sparse subspace are compact. Nevertheless, a purely geometric characterization of $\mathrm{ONL}_{\mathcal{P}_{\mathrm{Fin}}}$ remains an open question. In this paper, we employ higher index theory (see e.g. Lemma 4.4) to verify this property for $\ell^{p}$ coarsely embeddable metric spaces.
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into an $\ell^{p}$-space, then $X$ has operator norm localization for equi-approximable finite-rank projections.
Finally, we extend our results from the reduced setting to the maximal setting. By combining our main theorem with the techniques in Willett and Yu [2020]; Wang et al. [2024], we deduce the following result regarding the maximal coarse Baum-Connes conjecture:
@@
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ between the maximal and reduced Roe algebras induces an isomorphism in $K$-theory, i.e.,
@@
    For any subspace $Y$ in $X$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ between the maximal and reduced quotient Roe algebras induces an isomorphism in the level of $K$-theory, i.e.,
@@
    For any subspace $Y$ in $X$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
Our results motivate the following open questions:
@@
Let $X$ be a metric space with bounded geometry.
  • 1.
    If $X$ admits a coarse embedding into an $\ell^{p}$-space, is the coarse groupoid $G(X)$ necessarily $K$-amenable? Strong evidence for this can be found in Corollary 1.4.
  • 2.
    In Deng and Guo [2024], a twisted version of the coarse Baum-Connes conjecture with coefficients is introduced. It is natural to ask whether the twisted coarse Baum-Connes conjecture with coefficients holds for $X$ if $X$ admits a coarse embedding into an $\ell^{p}$-spaces?
  • 3.
    In parallel with the result of Chen et al. [2013], does the maximal coarse Baum-Connes conjecture hold for $X$ if $X$ admits a fibred coarse embedding into an $\ell^{p}$-space?
The paper is organized as follows. In Section 2, we recall some preliminaries on coarse geometry and ideal structures of Roe algebras. In Section 3, we shall prove our main theorem in four steps. Finally, we discuss applications to the relative and maximal coarse Baum-Connes conjectures as well as the operator norm localization property for finite-rank projections in Section 4.
@@ 2.1. Coarse groupoid
where $B(x,R)=\{x^{\prime}\in X\mid d(x,x^{\prime})\leq R\}$. Denote $\beta X$ the Stone-Čech compactification of $X$, which is the set of all ultrafilters on $X$ endowed with a topology making it a compact Hausdorff space such that $C(\beta X)\cong\ell^{\infty}(X)$. The reader is referred to Goldbring [[2022] ©2022] for details of ultrafilters and Stone-Čech compactification. For each $R>0$, we define the $R$-diagonal of $X\times X$ to be
@@
When $R=0$, the set $\Delta_{0}$ is the diagonal in $X\times X$. A subset of an $R$-diagonal is called an entourage (or a controlled set). The set of all entourages, denoted by $\mathcal{E}_{d}$, is called the coarse structure of $X$ associated with the metric. The coarse groupoid of $X$, denoted by $G(X)$, is defined to be
@@
The space $\beta X$ is the unit space of $G(X)$. The projection onto the first (second, resp.) coordinate $r:X\times X\to X$, ($s:X\times X\to X$, resp.) extends continuously to the range map $r:G(X)\to\beta X$ (the source map $s:G(X)\to\beta X$, resp.). The reader is referred to Skandalis et al. [2002] for details of the coarse groupoid. A subset $U\subset\beta X$ is called invariant if $r(\gamma)\in U$ is equivalent to $s(\gamma)\in U$ for any $\gamma\in G(X)$. For a subset $Y\subseteq X$, we define
@@
It is proved in [Špakula and Willett, 2017, Appendix C] that for any $\gamma\in G(X)$ with $s(\gamma)=\alpha\in\beta X$, there exists a partial translation $t:D\to R$ in $X$ such that $\gamma=\lim_{x\to\alpha}(t(x),x)\in G(X)$.
For any map $f:X\to Y$, we shall still denote $f$ the extended continuous map $\beta X\to\beta Y$ with a little abuse of notation. For any subset $A\subseteq\beta X$, we denote $G(X)|^{A}_{A}=r^{-1}(A)\cap s^{-1}(A)$, which is subgroupoid. Let $Y\subseteq X$ be a subset. Then the canonical inclusion map $i:Y\to X$ extends to an inclusion $i:\beta Y\to\beta X$. Thus, one can identify $\beta Y$ to the closure of $Y$ in $\beta X$. As a result, we have that $G(Y)=G(X)|^{\beta Y}_{\beta Y}$. For any invariant subset $U\subseteq\beta X$, we shall denote $U^{\beta Y}=\beta Y\cap U$, and it is direct to see that $U^{\beta Y}$ is $G(Y)$-invariant. A coarse ideal pair $(X,U)$ includes a metric space $X$ and an invariant subset $U$. Two pairs $(X,U)$ and $(Y,V)$ are coarsely equivalent if there exists a coarse equivalence $f:X\to Y$ such that $V$ is the smallest invariant open set containing $f(U)$. In this case, it is equivalent to say that $V$ is the minimal invariant open subset such that $U=f^{-1}(V)$.